# Berezinskii-Kosterlitz-Thouless transition in the XY model on the honeycomb lattice: A comprehensive Monte Carlo analysis

### v3

@@ Abstract
In this paper, we examined the Berezinskii-Kosterlitz-Thouless (BKT) phase transition in the two-dimensional XY model on the honeycomb lattice. To address its thermodynamical behavior, we conducted standard numerical Monte Carlo simulations with the simulated annealing protocol and entropic simulations based on the Wang-Landau algorithm. The inverse transition temperature ($\beta_{BKT}$) was determined using the second ($\Upsilon$) and fourth-order ($\Upsilon_{4}$) helicity modulus as the order parameter. We also used the fourth-order Binder cumulant of the magnetization as an additional tool to address the critical properties of the model. Our finite-size scaling analysis suggests $\beta_{BKT}=1.687(3)$ from $\Upsilon$ and $\beta_{BKT}=1.635(11)$ from $\Upsilon_{4}$. From the Binder cumulant crossing we found $\beta_{BKT}=1.724(2)$. These values deviate significantly from the expected theoretical value of $\sqrt{2}$. We believe that this discrepancy suggests that the theoretical assumptions regarding the analytical calculation may need to be revisited. Additionally, we calculated the vortex density and the formation energy of the vortex-antivortex pairs, where the obtained vortex formation energy is $2\mu=4.0(5)$. Upon comparison with the square lattice, our results support the notion of instability of the honeycomb lattice to support the spin long-range order and give additional backing to the critical behavior we found.
@@ I Introduction
In 2016, the Nobel Prize in Physics was awarded to David J. Thouless, F. Duncan M. Haldane, and J. Michael Kosterlitz for their groundbreaking contributions to the understanding of topological phase transitions and phases of matter, which have significantly advanced condensed matter physics. They demonstrated that, at sufficiently low temperatures, the magnetic moments of a two-dimensional lattice of atoms form vortex-antivortex pairs due to the presence of topological defects. However, a topological phase transition occurs when the system temperature increases. At a certain critical temperature ($\beta_{BKT}$), these pairs dissociate and roam freely through the lattice. This transition is called Kosterlitz-Thouless (KT) transition or BKT, in honor of Vadim Berezinskii, a deceased theoretical physicist who presented ideas similar to Kosterlitz and ThoulessBerezinskii (1972).
@@
The present study is motivated by two key factors. First, the BKT transition temperature for the square lattice is well established in the literature, with numerical calculations yielding $0.8935(1)$Hsieh et al. (2013). For the honeycomb lattice, some analytical studies predict an exact value of $1/\sqrt{2}\approx 0.707$Nienhuis (1982); Campostrini et al. (1996), which is confirmed by later calculations using cluster simulationsDeng et al. (2007) and diffusion map methodWang et al. (2021). Nonetheless, a detailed Monte Carlo (MC) study of the critical behavior on the honeycomb lattice appears to be absent in the existing literature. The second motivation stems from the ongoing debate regarding the appropriate parameter to characterize the BKT transition, given the limitations of computer simulations due to the finite size of simulated systems. The modulus of second-order helicity, which measures the system response to a global twist of spins in a given direction, is currently accepted as the parameter to describe this transitionMinnhagen and Kim (2003). The controversy arises because the discontinuity in the helicity modulus is not strongly dependent on the system size. While some researchers have shown that logarithmic corrections applied to the helicity modulus can yield accurate estimates of $\beta_{BKT}$sieh et al. (2013), it is also possible to use fourth-order helicity as an order parameter. This alternative presents a discontinuity at the critical temperature, potentially providing a more precise description of the BKT transition via entropic numerical simulations, where a high resolution search for the critical points is possibleCaparica (2014).
We primarily use the MC method and the Simulated Annealing (SA) protocolLandau and Binder (2014) to calculate the second- and fourth-order helicity. This helps us to estimate the inverse critical temperature $\beta_{BKT}$. Additionally, we use entropic simulations based on the Wang-Landau (WL) algorithmWang and Landau (2001a) to validate the SA results. The WL technique allows for a more precise determination of the critical region without being affected by critical slowdown. However, it is limited to small system sizes for models with continuous symmetry. Moreover, we will employ the finite-size scaling technique with logarithmic correction proposed by Y.D. Hsieh et al.Hsieh et al. (2013). We also calculate the vortex density and the energy formation of vortex pairs, which may provide insight into the stability of the honeycomb lattice compared to the square lattice and the location of the critical region.
@@ III.4 Thermodynamic quantities
$C_{v}=k_{B}\beta^{2}(\langle E^{2}\rangle-\langle E\rangle^{2}).$ (10)
$\chi=\beta(\langle M^{2}\rangle-\langle M\rangle^{2}).$ (11)
The Binder cumulant of the order parameter is
$B_{M}=1-\frac{\langle m^{4}\rangle}{3\langle m^{2}\rangle^{2}},$ (12)
where the angle brackets again denote a thermal average. This quantity is useful to determine the critical temperature since there is a crossing for different lattice sizes at a “fixed point” which is independent of the system size. Also, as $L\to\infty$, $B_{M}\to 1/3$ ($B_{M}\to 2/3$) after (before) the phase transition.
@@
$\lim_{L\to\infty}\Upsilon(\beta_{BKT})=\frac{2}{\pi\beta_{BKT}},$ (21)
This equality holds in the thermodynamic limit, as the system size draw near to infinity. The procedure involves calculating the helicity modulus for each temperature across various lattice sizes and comparing the results with the curve $r=2/(\beta\pi)$. The intersection points determine the critical temperature for each lattice size. To estimate the BKT transition temperature using the fourth-order helicity modulus (Eq. 20), one simply identifies the temperature at which the function reaches its minimum.
@@
Figure 2: (top) Specific heat with average energy in the inset, (middle) magnetic susceptibility with average magnetization in the inset and (bottom) Binder cumulant versus temperature for different lattices sizes. The dots correspond to the SA results and the full line to the WL ones.
@@ IV Results
All simulations were conducted on lattices with sizes ranging from $L=6$ to $L=128$ with periodic boundary conditions. For the SA protocol, we set the initial temperature to $\beta_{0}=0.1$ and the final temperature to $\beta_{n}=3$, with $n=100$ and $m_{s}=10^{6}$. This involved a total of $10^{8}$ MC sweeps for each temperature. Additionally, ten independent simulations were performed for each lattice size to compute statistical errors. For the largest lattice size, all spin configurations were saved at each temperature to analyze the structural order of the spins and calculate the vortex density.
Table 1: Critical temperatures obtained via FSS of the specific heat and susceptibility peaks from both SA and WL techniques.
SA WL
$C_{v}$ 0.667(12) 0.671(3)
$\chi$ 0.642(21) 0.639(2)
@@
Figure 2 presents various thermodynamic quantities as a function of temperature for both SA (indicated by dots) and WL (represented by solid lines) procedures. Since the WL simulation is constrained to smaller system sizes, we present results only for $L$ values up to 32. As shown, the specific heat approaches $1/2$ as $T\to 0$, regardless the lattice size, which aligns with the expectations of the equipartition theorem. Although both the specific heat and susceptibility exhibit a peak slightly above $T=0.6$, these peaks are not used to determine the BKT transition temperature due to their lack of accuracy. Additionally, the peak of the specific heat does not display a consistent behavior as the system size increases. However, a FSS analysis of the critical temperature using this data yields values that are lower than the expected theoretical value of $1/\sqrt{2}$, as summarized in Table 1.
Figure 3: Average second-order helicity as a function of inverse temperature for different lattice sizes obtained by SA. The theoretical solid curve $2/\pi\beta$ indicates the critical temperature through the intersection with the helicity curve.
Some authors use the Binder cumulant of the order parameter crossing technique together with another sophisticated numerical procedure to determine the critical temperatureSilva et al. (2022). Figure 2 also presents the behavior of this quantity as a function of temperature for several lattice sizes. As expected, the Binder cumulants approach a value of $1/3$ in the disordered (high-temperature) state. Although for the ordered (low temperature) state they approach the value of $2/3$.
Figure 4: Average fourth-order helicity modulus as a function of the inverse temperature for various lattice sizes.
For a more precise determination of the critical temperature, we use the helicity modulus obtained from the SA procedure as a function of the inverse temperature for various lattice sizes, as illustrated in Fig. 3. We use inverse temperature in this context because the SA results depend on $\beta$. This choice also minimizes rounding errors during the finite-size scaling analysis. The theoretical curve, represented by $2/\pi\beta$, is also shown. The discontinuity of the helicity modulus near the transition is not very pronounced. As a discontinuity occurs in the thermodynamic limit, we can use the intersection of the theoretical curve $2/\pi\beta$ with the simulated values to estimate the critical temperature. As the lattice size increases, the intersection point shifts to the right, converging toward the critical temperature in the thermodynamic limit. A key challenge lies in accurately identifying this intersection, as the SA data are discrete. Moreover, obtaining a larger sample size with SA would incur significant computational costs. This issue does not arise with WL simulations, but they are limited as system size increases.
Figure 5: Evolution of transition temperatures with system size obtained from both $\langle\Upsilon\rangle$ and $\langle\Upsilon_{4}\rangle$.
@@
Based on this, we individually calculated the values of $\beta_{BKT}$ for each lattice and performed a finite-size scaling analysis, as shown in Fig. 5, for both $\langle\Upsilon\rangle$ and $\langle\Upsilon_{4}\rangle$. Using the helicity modulus, we obtained $\beta_{BKT}=1.687\pm 0.003$. It is important to note that this value is significantly higher than the theoretical prediction of $\sqrt{2}$. Although the WL technique yields $\beta_{BKT}=1.627\pm 0.004$. The difference is due to the maximum limited size of $L=32$ used in the latter. But this value is still higher than the theoretical one.
@@
Obtaining the critical temperature through $\langle\Upsilon_{4}\rangle$ is relatively straightforward. We calculated $\beta_{c}$ for each lattice size and performed the FSS to determine $\beta_{BKT}$, as shown in Fig. 5. We found $\beta_{BKT}=1.635\pm 0.011$ using SA and $\beta_{BKT}=1.648\pm 0.007$ using the WL method. There is a slight discrepancy between the results from the two approaches because the WL method allows a higher-resolution search for the critical temperature corresponding to $\langle\Upsilon_{4}\rangle$. Additionally, it is important to note that the values obtained from $\langle\Upsilon\rangle$ and $\langle\Upsilon_{4}\rangle$ do not agree even within their error margins. This discrepancy may be due to the gradual approach of the minimum of $\langle\Upsilon_{4}\rangle$ towards the critical temperature in the thermodynamic limit, suggesting that the scaling relation given by Eq. 22 may not be appropriate in this context. Nevertheless, these results indicate that the theoretical predictions are not supported by numerical analysis.
Figure 6: (top) Crossing of the Binder cumulant for several lattice sizes. Some lattice sizes were omitted for clarity. (bottom) Inverse temperature of the crossings of the Binder cumulant as a function of different lattice sizes. The data was obtained via SA considering $L=6,8,12,16,$ and $24$ as the smallest lattice.
An alternative yet powerful approach for identifying critical properties and determining the transition temperature is the analysis of the crossing points of the Binder cumulant, a technique widely used in FSS studiesBinder (1981). Figure 6 illustrates the cumulant crossings for different system sizes as a function of the inverse temperature obtained through SA. Despite the apparent randomness of the crossings, it is evident that the cumulant exhibits scaling behavior. Albuquerque et al. employed a sophisticated extrapolation scheme of the crossings to obtain the critical temperature of a modified XY model on the square latticeSilva et al. (2022). However, since we are primarily interested in the general aspects of the cumulant crossings to support our findings, we will use a method similar to that described in Ref. Liu et al. (2018). Figure 6 (bottom) also shows how the inverse temperature at which the crossings occur varies with lattice sizes. A rough estimate of the inverse temperature in the thermodynamic limit gives $\beta_{cross}=1.724\pm 0.002$. Notably, this value is even higher than those obtained from the helicity modulus compared to the theoretical prediction. Very recently, F. Jiang reported neural network calculations of the same problemJiang (2024) and found a value of $\approx 1.75$ for $\beta_{BKT}$, which further supports our findings.
@@
It is helpful to illustrate the topology of spin configurations at different temperatures as the BKT transition passes through. Figure 7 displays six configurations for the lattice size of $L=128$. The color pattern indicates the individual orientation of each spin and follows the following scale (in radians): dark red – $+\pi/2$ rad; dark blue – $-\pi/2$ rad; white – 00 rad; black – ±$\pm 2\pi$ rad. For the temperature $T=0.25$, below the critical temperature, we can see the formation of a pattern in the orientation of the spins (predominance of spins oriented “vertically”) but it is not uniform throughout the lattice; this occurs due to the presence of vortex-antivortex pairs. For $T=0.57$, in the region where the transition occurs, it is easy to see how the system becomes quite heterogeneous, possibly due to the dissociation of the vortex-antivortex pairs. As the temperature rises above $T_{BKT}$, the spins become misaligned, leading to an increase in disorder until complete dissociation occurs. At this point, the system reaches a saturation point in the number of vortices.
Figure 7: Spin configurations given by a color gradient of the spin orientation angle in radians for lattice size $L=128$ and some values of temperatures: (a) $T=0.25$, (b) $T=0.57$, (c) $T=0.69$, (d) $T=0.76$ , (e) $T=1.19$ and (f) $T=25$. The dark red color indicates spin “up” ($+\pi/2$) and dark blue spin “down” ($-\pi/2$).
@@
where $2\mu$ is the angular coefficient of this function. Using this relationship, we adjusted the vortex density values close to the critical temperature and obtained $2\mu=4.0\pm 0.5$ in units of $J$. This value is less than half of what was obtained for the square latticeGupta and Baillie (1992). This result is justified by the idea of the instability of the honeycomb lattice to support the spin long-range order, as investigated by Wojtkiewicz et al.Wojtkiewicz et al. (2023).
Figure 8: Vortex density as a function of temperature for $L=128$.

### v4

@@ Abstract
In this paper, we thoroughly examined the Berezinskii-Kosterlitz-Thouless (BKT) phase transition in the two-dimensional XY model on the honeycomb lattice. To address its thermodynamical behavior, we combined standard numerical Monte Carlo simulations with the simulated annealing (SA) protocol and entropic simulations based on the Wang-Landau (WL) algorithm. The transition temperature was determined using the second ($\Upsilon$) and fourth-order ($\Upsilon_{4}$) helicity modulus as the order parameter. Our best finite-size scaling analysis suggests $T_{BKT}=0.575(8)$ from SA and $T_{BKT}=0.576(3)$ from WL. These values deviate significantly from the expected theoretical value of $1/\sqrt{2}$. We believe that this discrepancy suggests that the theoretical assumptions regarding the analytical calculation may need to be revisited. Additionally, we calculated the vortex density and the formation energy of the vortex-antivortex pairs, where the obtained vortex formation energy is $2\mu=5.80(12)$. Upon comparison with the square lattice, our results support the notion of instability of the honeycomb lattice to support the spin long-range order and give additional backing to the critical behavior we found.
@@ I Introduction
In 2016, the Nobel Prize in Physics was awarded to David J. Thouless, F. Duncan M. Haldane, and J. Michael Kosterlitz for their groundbreaking contributions to the understanding of topological phase transitions and phases of matter, which have significantly advanced condensed matter physics. They demonstrated that, at sufficiently low temperatures, the magnetic moments of a two-dimensional lattice of atoms form vortex-antivortex pairs due to the presence of topological defects. However, a topological phase transition occurs when the system temperature increases. At a certain critical temperature ($T_{BKT}$), these pairs dissociate and roam freely through the lattice. This transition is called Kosterlitz-Thouless (KT) transition or BKT, in honor of Vadim Berezinskii, a deceased theoretical physicist who presented ideas similar to Kosterlitz and ThoulessBerezinskii (1972).
@@
The present study is motivated by two key factors. First, the BKT transition temperature for the square lattice is well established in the literature, with numerical calculations yielding $0.8935(1)$Hsieh et al. (2013). For the honeycomb lattice, some analytical studies predict an exact value of $1/\sqrt{2}\approx 0.707$Nienhuis (1982); Campostrini et al. (1996), which is confirmed by later calculations using cluster simulationsDeng et al. (2007) and diffusion map methodWang et al. (2021). Nonetheless, a detailed Monte Carlo (MC) study of the critical behavior on the honeycomb lattice appears to be absent in the existing literature. The second motivation stems from the ongoing debate regarding the appropriate parameter to characterize the BKT transition, given the limitations of computer simulations due to the finite size of simulated systems. The modulus of second-order helicity, which measures the system response to a global twist of spins in a given direction, is currently accepted as the parameter to describe this transitionMinnhagen and Kim (2003). The controversy arises because the discontinuity in the helicity modulus is not strongly dependent on the system size. While some researchers have shown that logarithmic corrections applied to the helicity modulus can yield accurate estimates of $T_{BKT}$sieh et al. (2013), it is also possible to use fourth-order helicity as an order parameter. This alternative presents a discontinuity at the critical temperature, potentially providing a more precise description of the BKT transition via entropic numerical simulations, where a high resolution search for the critical points is possibleCaparica (2014).
We primarily use the MC method and the Simulated Annealing (SA) protocolLandau and Binder (2014) to calculate the second- and fourth-order helicity. This helps us to estimate the inverse critical temperature $T_{BKT}$. Additionally, we use entropic simulations based on the Wang-Landau (WL) algorithmWang and Landau (2001a) to validate the SA results. The WL technique allows for a more precise determination of the critical region without being affected by critical slowdown. However, it is limited to small system sizes for models with continuous symmetry. Moreover, we will employ the finite-size scaling technique with logarithmic correction proposed by Y.D. Hsieh et al.Hsieh et al. (2013). We also calculate the vortex density and the energy formation of vortex pairs, which may provide insight into the stability of the honeycomb lattice compared to the square lattice and the location of the critical region.
@@ III.4 Thermodynamic quantities
$C_{v}=\frac{1}{k_{B}T^{2}}(\langle E^{2}\rangle-\langle E\rangle^{2}).$ (10)
$\chi=\frac{1}{k_{B}T}(\langle M^{2}\rangle-\langle M\rangle^{2}).$ (11)
@@
It is important to state that equations 16 and 19 must be multiplied by the factor of $\dfrac{4}{3\sqrt{3}}$ in the case of the honeycomb lattice. This account for the ratio of spin densities when compared to the square lattice.
@@
$\lim_{L\to\infty}\Upsilon(\beta_{BKT})=\frac{2T_{BKT}}{\pi},$ (20)
This equality holds in the thermodynamic limit, as the system size draw near to infinity. The procedure involves calculating the helicity modulus for each temperature across various lattice sizes and comparing the results with the curve $r=2T/\pi$. The intersection points determine the critical temperature for each lattice size. To estimate the BKT transition temperature using the fourth-order helicity modulus (Eq. 19), one simply identifies the temperature at which the function reaches its minimum.
@@
Figure 2: (top) Specific heat with its maximum value as a function of $L$ in the inset and (bottom) magnetic susceptibility with average magnetization in the inset for different lattices sizes. The dots correspond to the SA results and the full line to the WL ones. The vertical dashed line indicates the theoretical expected value for $T_{c}$.
@@ IV Results
All SA simulations were conducted on lattices with sizes ranging from $L=6$ to $L=128$ with periodic boundary conditions. For the SA protocol, we set the initial temperature to $\beta_{0}=0.1$ and the final temperature to $\beta_{n}=2$, with an increment of $\Delta\beta=0.01$. This involved a total of $10^{5}$ MC sweeps for thermalization part and the same amount for acquire the thermal averages. Additionally, 20 independent simulations were performed for each lattice size to compute statistical errors via the bootstrap method. For the largest lattice size, all spin configurations were saved at each temperature to analyze the structural order of the spins and calculate the vortex density.
Figure 3: Average second-order helicity as a function of inverse temperature for different lattice sizes obtained by SA. The inset shows the same result obtained by WL. The theoretical solid curve $2T/\pi$ indicates the critical temperature through the intersection with the helicity curve. The vertical dashed line indicates the theoretical expected value for $T_{c}$.
@@
Figure 2 presents various thermodynamic quantities as a function of temperature for both SA (indicated by dots) and WL (represented by solid lines) procedures. Since the WL simulation is constrained to smaller system sizes, we present results only for $L$ values up to $48$. It is worth-noting the high resolution of the WL results when compared to the SA ones. The specific heat value approaches $1/2$ as $T\to 0$, regardless the lattice size, which aligns with the expectations of the equipartition theorem. Although both the specific heat and susceptibility exhibit a peak slightly above $T=0.6$, these peaks are not used to determine the BKT transition temperature due to their lack of accuracy. Additionally, the peak of the specific heat does not display a consistent behavior as the system size increases (as can be seen in the inset of the same figure).
Figure 4: Average fourth-order helicity modulus as a function of the inverse temperature for various lattice sizes. The vertical dashed line indicates the theoretical expected value for $T_{c}$.
For a precise determination of the critical temperature, we use the helicity modulus obtained from both SA and WL procedures as a function of the temperature for various lattice sizes. Figure 3 shows some of these results for SA and the inset shows the results for WL up to lattice $L=48$. The theoretical curve, represented by $2T/\pi$, is also shown. The discontinuity of the helicity modulus near the transition is not very pronounced. As a discontinuity occurs in the thermodynamic limit, we can use the intersection of the theoretical curve $2T/\pi$ with the simulated values to estimate the critical temperature. As the lattice size increases, the intersection point shifts to the right, converging toward the critical temperature in the thermodynamic limit. A key challenge lies in accurately identifying this intersection, as the SA data are discrete. Moreover, obtaining a larger sample size with SA would incur significant computational costs. This issue does not arise with WL simulations, but they are limited as system size increases.
Figure 5: Finite-size scaling for both $\langle\Upsilon\rangle$ (top) and $\langle\Upsilon_{4}\rangle$ (bottom) obtained via SA and WL approaches.
@@
Based on this, we individually calculated the values of $T_{c}$ for each lattice and performed a finite-size scaling analysis using an improved method proposed by Hsieh et al.Hsieh et al. (2013). We used one of their forms for the scaling relation, i.e.
$T_{c}(L)=T_{BKT}(\infty)+\frac{a}{\ln^{2}(bL)},$ (22)
for all simulated $T_{c}$. Figure 5 (top) shows our FSS results for both $\langle\Upsilon\rangle$ and $\langle\Upsilon_{4}\rangle$. Using the helicity modulus, we obtained $T_{BKT}=0.575(8)$ via SA and $T_{BKT}=0.576(3)$ via WL. It is important to note that this value is significantly lower than the theoretical prediction of $1/\sqrt{2}$. Although the WL technique is limited to the system size, the result is highly accurate. This is not a surprise since the logarithmic correction proposed by Weber and MinnhagenWeber and Minnhagen (1988) and improved by Hsieh et al. to the study of the BKT transition is supposed to give accurate results. This agrees with the empirical observation that Equation 16 is valid to extremely good approximation even for small lattice sizes. Very recently, F. Jiang reported neural network (NN) calculations of the same problemJiang (2024) and found a value of $0.572(3)$ for $T_{BKT}$, which is in very good agreement with our findings.
@@
Obtaining the critical temperature through $\langle\Upsilon_{4}\rangle$ is relatively straightforward. We calculated $T_{c}$ for each lattice size and performed the FSS to determine $T_{BKT}$, as shown in Fig. 5 (bottom). We found $T_{BKT}=0.551(11)$ using the SA method, and $T_{BKT}=0.568(1)$ from the WL. It is important to note that the values obtained from $\langle\Upsilon\rangle$ and $\langle\Upsilon_{4}\rangle$ using SA do not agree within their error margins. This discrepancy may arise from the gradual approach of the minimum of $\langle\Upsilon_{4}\rangle$ toward the critical temperature in the thermodynamic limit, suggesting that the scaling relation presented in Eq. 22 may not be appropriate in this context. Another challenge is that the data obtained via SA near the minimum exhibit significant fluctuations, which can lead to an inaccurate determination of $T_{c}$. In contrast, the data from the WL method allow for a higher-resolution search for the corresponding critical temperature. However, there is still a slight discrepancy when compared to the $T_{BKT}$ obtained using the helicity modulus but agrees within the error bar with the NN calculations. The FSS analysis using $\langle\Upsilon_{4}\rangle$ may require all logarithmic corrections to be taken into account in a more precise investigation of the BKT transition. Nevertheless, our results still suggest that the theoretical predictions are not fully supported by the numerical analysis. The close form of $T_{c}$ for $O(N)$ vector models proposed by NienhuisNienhuis (1982) may not capture all the universal picture of this kind of model.
Figure 6: Spin configurations for lattice size $L=128$ and some values of temperatures: (a) $T=0.500$, (b) $T=0.549$, (c) $T=0.575$, (d) $T=0.60$ , (e) $T=0.71$ and (f) $T=0.91$. The red stars (blue dots) indicate the location of a vortex (antivortex).
@@
It is helpful to illustrate the topology of spin configurations at different temperatures as the BKT transition passes through. Figure 6 displays six configurations for the lattice size of $L=128$. The red stars (blue dots) indicate the location of a vortex (antivortex). For the temperature $T=0.50$, below the critical temperature, one can see some vortex-antivortex pairs bound. The vortex density is very low, ideally tending to zero in the thermodynamic limit. The density rises very slowly with $T$. For $T\approx 0.57$, in the region where the transition occurs, there is a sudden increase in vortex density as vortex-antivortex pairs unbind. This is not a sharp jump (as in first-order phase transitions), but a rapid crossover. As the temperature rises above $T_{BKT}$, vortices and antivortices proliferate, leading to an increase in disorder until complete dissociation occurs. At this point, the system reaches a saturation point in the number of vortices.
@@
where $2\mu$ is the angular coefficient of this function. Using this relationship, we adjusted the vortex density values close to the critical temperature and obtained $2\mu=5.80(12)$ in units of $J$. This value is lower than the one obtained for the square latticeGupta and Baillie (1992). This result is justified by the idea of the instability of the honeycomb lattice to support the spin long-range order, as investigated by Wojtkiewicz et al.Wojtkiewicz et al. (2023).
Figure 7: Vortex density as a function of temperature for $L=128$. The inset shows the linear fitting given by Eq. 24.